Copyright (C) 1993 Mark R. Rinfret, MRsoftware All Rights Reserved.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

MRIconSort

Amiga WorkBench Icon Sorter Version 1.0.0 March 1993

Mark R. Rinfret

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 Mark R. Rinfret, MRsoftware

1 Permissions

     Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 Mark R. Rinfret, MRsoftware
     All Rights Reserved.

MRIconSort is freely distributable shareware. It may be redistributed via magnetic media or electronic bulletin board systems (BBS). It is unlawful to charge a fee for this software other than reasonable distribution/media fees.

The author assumes no responsibilty for damages incurred as a result of using MRIconSort. USE AT YOUR OWN RISK! No warranties, either expressed or implied, are made with regard to its use on any system.

Once you have tried MRIconSort and have decided to continue using it, you are requested to submit five dollars ($5.00 U.S.) to the author. You should also provide the following information:

         Your mail address
         Your e-mail address(es), if any
         The version of MRIconSort you are registering

Send this to:

         MRsoftware
         348 Indian Avenue
         Portsmouth, RI  02871

2 MRIconSort User’s Manual

2 Overview

MRIconSort (Mark Rinfret’s Icon Sort - go ahead, call it "Mister Icon Sort" :-) is a tool which will alphabetize and align your WorkBench icons, creating order where there is chaos.

Though the WorkBench interface has its niceties, things can run amuck when you have drawers containing many icons. Things become especially confusing when most of those icons look the same (large groups of pictures, drawers, etc.). The same visual cues that make the WorkBench powerful for drawers containing a few icons render it useless when you have many similarly sized and shaped icons.

MRIconSort scans a directory (or a full directory hierarchy) and rearranges the icons to make navigating a WorkBench window much more orderly. You can choose to group icons by files first, directories first or mixed (no preference). Then, icons are alphabetized and aligned from left to right. The vertical spacing of the icons is arranged such that the tallest icon in any row determines the height of that row and all icon text sits on the same baseline, making it easy to scan all names without getting a case of the "jumpy eyeball".

MRIconSort knows about icons that have been left out on the desktop (it reads the :.backdrop file) and leaves them alone. You have your choice of right "margin": you can either use the WorkBench maximum or MRIconSort will attempt to get the "snapshot" coordinates of the window associated with each drawer.

MRIconSort can also be told to create default icons for files or directories which don’t have icons.

MRIconSort requires AmigaDOS 2.04 and up. It will run from either the Shell or the WorkBench.

MRIconSort’s default behavior can be customized through the definition of a preferences file named ‘MRIconSort.prefs’ (see Section 3.3 [Installing MRIconSort], page 6).

4 MRIconSort User’s Manual

3 Installing MRIconSort

3.1 Unpacking the MRIconSort Archive

You most likely received MRIconSort as an Lha archive. It will unpack into its own directory hierarchy with the following command:

     lha -xa x MRIconSort.lha

The actual archive name may vary from ‘MRIconSort.lha’ since it may include a version number. If the archive is unpacked correctly, a new directory will be created containing the MRIconSort program, documentation and environment variable directories.

3.2 Installing the Software

Make the MRIconSort directory the current directory by using the CD command.

You can install the MRIconSort program in any suitable directory which is a part of your command search path (i.e. defined by the AmigaDOS PATH command). For example, if you have a directory named ‘SYS:Tools’, you could copy the MRIconSort program there with

     COPY CLONE MRIconSort MRIconSort.info SYS:Tools

It is assumed that you have at least one directory on your system where document files are stored (though this is not a requirement). It is quite likely that you have several documentation directories (quite likely not sharing a common hierarchy).

The documentation for MRIconSort comes in three forms:

• A plain text file named ‘MRIconSort.doc’

• An AmigaGuide(tm) file named ‘MRIconSort.guide’

• A TeX ‘DVI’ file named ‘MRIconSort.dvi’.

Copy any or all of these to the appropriate online documentation area.

3.3 Edit MRIconSort Preferences

When MRIconSort starts up, it first looks for a file name ‘MRIconSort.prefs’ in the current directory. If this file doesn’t exist, it then looks for the file ‘ENV:MRIconSort.prefs’. The default

‘MRIconSort.prefs’ file is stored in the MRIconSort installation directory as ‘Env/MRIconSort.prefs’.

This file contains MRIconSort configuration parameters which customize MRIconSort behavior to your unique requirements.

The supplied preferences file is self-documenting. The ‘MRIconSort.prefs’ file is a text file organized as one directive per line with lines of comments (lines beginning with a semi-colon) interspersed with lines of directives. Here is the supplied ‘MRIconSort.prefs’ file:

; This is the MRIconSort configuration file (environment variable).
; All directives are included here. Those preceded by ";" are ignored.
; To activate any directive, just edit this file to remove the
; leading semi-colon from that line.
;
; All - Process all levels (recursive directory scan).
;All
;
; FilesFirst    - Place all file icons before directory icons.
FilesFirst
;
; DirsFirst     - Place all directory icons before file icons.
;DIRSFIRST
;
; Mixed         - Mix file and directory icons on the same line.
;Mixed
;
; XSpace = <number> - Set horizontal spacing between icons (default = 8).
XSpace = 8
;
; YSpace = <number> - Set vertical spacing between icons (default = 2).
YSpace = 2
;
; MaxWidth -    - Use the maximum WorkBench screen width.
;                 (If absent, use icon's window snapshot.)
;MaxWidth

Once these changes are made, copy the environment variable directory to both the ENV: and ENVARC: directories:

     COPY CLONE ALL Env ENV:
     COPY CLONE ALL Env ENVARC:

If you are make changes to your global MRIconSort Preferences, remember to copy the ENV:MRIconSort.prefs file to ENVARC:. Otherwise, your changes will be lost the next time you reboot your system.

8 MRIconSort User’s Manual Chapter 4: Using MRIconSort 9

4 Using MRIconSort

MRIconSort can be invoked from the WorkBench or the Shell. There are distinct differences in its behavior under each environment.

10 MRIconSort User’s Manual

5 Shell Operation

When run from the Shell, MRIconSort attempts to read the preferences file, ‘MRIconSort.prefs’, first from the current directory and, if no file is found there, from ‘ENV:MRIconSort.prefs’. Any command line parameters then override the preferences settings. The basic format of the command line is:

 MRIconSort <directory_name> <options>

The MRIconSort command template is:

 Dir/A,DirsFirst/S,FilesFirst/S,Mixed/S,Create/S,All/S,MaxWidth/S,
 XSpace/K/N,YSpace/K/N

MRIconSort’s parameters are defined as follows:

• Dir This is the name of the directory whose icon’s are to be sorted/created.

• DirsFirst This is a switch which indicates that directory icons are to appear before file icons. The default is FilesFirst.

• FilesFirst This is a switch which indicates that file icons are to appear before directory icons. This is the default behavior.

• Mixed This is a switch which indictes that no preference is to be given to the order of file and directory icons (i.e. they will be mixed).

• Create This is a switch which turns on icon creation for any files or directories not having an icon.

• All This switch instructs MRIconSort to sort all levels (descend the directory tree).

• MaxWidth Normally, MRIconSort will attempt to obtain the WorkBench window dimension "snapshot" for the drawer being organized and use the window width as the right-hand margin for each row of icons. However, if this switch is specified, MRIconSort will use the current WorkBench screen width.

• XSpace = <pixels> This parameter specifies the horizontal spacing factor. MRIconSort defaults to 8 pixels.

• XYspace = <pixels> This parameter specifies the vertical spacing factor. MRIconSort defaults to 2 pixels.

6 WorkBench Operation

When MRIconSort is invoked from the WorkBench, it will first attempt to read its preferences settings from the ‘ENV:MRIconSort.prefs’ file and then will override these settings with any ToolType entries it may find in its tool icon. The MRIconSort ToolType entries are identical in name and syntax to the command line template definitions (see Chapter 5 [Shell Operation], page 11) with the exception of the Dir parameter which cannot be specified as a ToolType.

Multiple directories can be processed when running from the WorkBench. Simply shift-select the icons for the drawers you want to process.

14 MRIconSort User’s Manual

7 Things to Do

I intend to evolve and improve MRIconSort over time. Your input will be greatly appreciated. Some of the changes planned (not necessarily in any particular order) are:

• Grid Snap This option will force icons to snap to an invisible grid.

• Window Dimension Setting This option will reset all drawer window dimensions to a user-specified value.

• Icon Creation Exclusion List This feature will inhibit icon creation for filenames matching a list of patterns (e.g. *.c, *.o, etc.).

16 MRIconSort User’s Manual Concept Index 17

Concept Index

E

Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

I

Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

M

MRIconSort Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

O

Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

P

Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Plans for the Future . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

S

Shell Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

U

Using MRIconSort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

W

WorkBench Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

18 MRIconSort User’s Manual

Short Contents

1 Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Installing MRIconSort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Using MRIconSort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 Shell Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6 WorkBench Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7 Things to Do . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

ii MRIconSort User’s Manual

Table of Contents

1 Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Installing MRIconSort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Unpacking the MRIconSort Archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Installing the Software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Edit MRIconSort Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 Using MRIconSort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 Shell Operation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6 WorkBench Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7 Things to Do . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

iv MRIconSort User’s Manual